What is the predetermined overhead rate, and why do companies use this calculation to apply manufacturing overhead? Write in simple words in one paragraph.
Answer #1

a predetermined overhead rate is an allocation rate,which is used to apply the budgeted manufacturing overheads to cost object.in simple words it is a rate which allocates the indirect manufacturing cost of a product .allocation basis are direct labour hours,machine hours etc.. . there are many indirect manufacturing cost incurred by a company in its production,the predetermined overhead rate helps the company to allocate the manufacturing overheads in setting the price of the product.

predetermined overhead rate=budgeted overhead/budgeted activity level

consider an example

ABC co produces a product xactive.direct material per unit is $12 and direct labour per unit is $20,the direct labour rate per hour is $5.the budgeted overhead for production is $50000 and it is allocated on the basis of direct labour hours.assume the company sells 500 units.determine the predetermined overhead?

answer:-

predetermined overhead rate=budgeted overhead/budgeted activity level

budgeted overhead=$50000

overheads are absorbed on the basis of direct labour hours so we need to find the total labour hours of 5000 units

direct labour hours for 1 unit=$20/$5=4 hours

direct labour hours for 500 units=4*500=2000 hours

this 1500 hours is the budgeted activity level.

predetermined overhead rate=$50000/2000=$25 per hour